OpenStudy (anonymous):

no solution, one solution, or infinitely many solution. x+y=3 x+y=-3

OpenStudy (anonymous):

1 solution I think

OpenStudy (anonymous):

help with mine please http://openstudy.com/study#/updates/531113b7e4b0cecd2345e473

OpenStudy (***[isuru]***):

u can take x + y = 3 ----(1) and x + y = -3 ---(2) then (1) - (2) this will leave u with 0 = 6 so.. u got no x or y values remaining.. which means there are no solutions to this simultaneous equations.. besides if u know about graphing... u can see these 2 r parallel lines ( since their slopes are equal) parallel lines never intersect each other... so.. It also concludes that there are no solutions to this equation :)